# Optimal Nuisance Function Tuning for Estimating a Doubly Robust Functional under Proportional Asymptotics

### v1

@@ Appendix G Additional details and results for the simulation study
    We illustrate the asymptotic variances of the debiased estimators across a range of values for the nuisance parameters when $c=0.5$. We then compare the optimal tuning parameters to those that minimize the prediction error (i.e., the prediction-optimal tuning parameters), as in the simulations presented in the main text. For the sake of completeness, we also present the results for $c=2$ where we do not trim the y-axis (unlike the main text).
@@
Throughout our simulations, we consider the case where the nuisance functions are tuned identically, i.e., $\lambda_{1}=\lambda_{2}=\lambda$. We consider a discrete grid of 100 values for $\lambda$ (ranging from 0.05 to 10). For each of the de-biased estimators, we find the value of $\lambda$ resulting in the lowest asymptotic variance for each estimator and compare it to the prediction-optimal $\lambda$ (i.e., $\lambda_{1}=\lambda_{2}=c$).
The code will be released at an anonymous GitHub link.
G.1 Simulations with two splits
G.1.1 Setup
@@ G.1.2 Results
G.2 Simulations with three splits
G.2.1 Setup
@@ G.2.2 Results
G.3 Verification of prediction-optimal tuning parameters

### v2

@@ Appendix H Additional Details and Results for the Simulation Study
    We illustrate the asymptotic variances of the debiased estimators across a range of values for the nuisance parameters when $c=0.5$ and $c=2$. We then compare the optimal tuning parameters (that minimize the variance) to those that minimize the prediction error (i.e., the prediction-optimal tuning parameters), as in the simulations presented in the main text. For $c=2$, where we do not trim the y-axis (unlike the main text).
@@
  4. 4.
    We illustrate that the parametric bootstrap approach performs well for estimating the asymptotic variance of the debiased estimators.
@@
Throughout our simulations, we consider the case where the nuisance functions are tuned identically, i.e., $\lambda_{1}=\lambda_{2}=\lambda$. We consider a discrete grid of 100 values for $\lambda$ (ranging from 0.05 to 10). For each of the debiased estimators, we find the value of $\lambda$ resulting in the lowest asymptotic variance for each estimator and compare it to the prediction-optimal $\lambda$ (i.e., $\lambda_{1}=\lambda_{2}=c$).
The code will be released at the GitHub link: https://github.com/zixiaowang17/Optimal-Nuisance-Function-Tuning.
H.1 Bias and Variance in Settings with Two Splits
H.1.1 Setup
@@ H.1.2 Results
H.2 Bias and Variance in Settings with Three Splits
H.2.1 Setup
@@ H.2.2 Results
H.3 Verification of Prediction-Optimal Tuning Parameters
@@
H.4 Parametric Bootstrap for Estimating the Asymptotic Variance
Our next set of results evaluate the performance of the parametric bootstrap approach described in Appendix G. In particular, we compared the estimated variance of the debiased ECC estimators (with $10,000$ bootstrap samples) to their true variances, obtained by Monte Carlo integration with $10,000$ samples.
H.4.1 Settings with Two Splits
Figure 12 compares the estimated variance of the ECC estimators with the true variance in the setting with two splits and $c=0.5$. Table 1 summarizes the results across $\lambda$ values. We find that the estimated variance was very close to the true variance for each $\lambda$ value.
Table 1: Ratio of the true standard error to the bootstrap estimated standard error of the debiased ECC estimators in the setting with two splits and $c=0.5$. The entries show summary statistics of the ratios calculated across the 100 different $\lambda$ values. Ratios greater than 1 indicate that the true standard error is greater than the bootstrap estimated standard error.
Estimator Min 1st Qu. Median Mean 3rd Qu. Max
INT (integral-based) 0.995 0.996 0.999 1.001 1.004 1.019
NR (Newey–Robins) 0.992 0.992 0.993 0.996 0.996 1.035
IF (doubly robust) 0.983 0.989 0.992 0.991 0.993 0.996
Figure 12: Bootstrap estimated (black dots) versus true (red line) variance of the debiased ECC estimators in the setting with two splits and $c=0.5$ across the $\lambda$ values.
Analogous results in the $c=2$ setting are given in Figures 13 and Table 2. As noted in the main text, the variance blows up when $\lambda\approx 1.5$ due to the bias correction constant diverging for the Integral-based estimator. We also noticed a similar phenomenon for the Newey and Robins estimator for $\lambda$ close to 0. We plan to explore these subtitles in detail in the future. Compared to the $c=0.5$ setting, the differences between the estimated and true variances were larger, especially for the integral-based estimator.
Figure 13: Bootstrap estimated (black dots) versus true (red line) variance of the debiased ECC estimators in the setting with two splits and $c=2$ across the $\lambda$ values.
Table 2: Ratio of the true standard error to the bootstrap estimated standard error of the debiased ECC estimators in the setting with two splits and $c=2$. The entries show summary statistics of the ratios calculated across the 100 different $\lambda$ values. Ratios greater than 1 indicate that the true standard error is greater than the bootstrap estimated standard error.
Estimator Min 1st Qu. Median Mean 3rd Qu. Max
INT (integral-based) 0.636 0.787 0.813 0.831 0.824 0.995
NR (Newey-Robins) 0.862 0.872 0.913 0.920 0.965 1.005
IF (doubly robust) 0.894 0.917 0.936 0.941 0.966 0.995
H.4.2 Settings with Three Splits
Figure 14 and Table 3 summarize the simulation results in the setting with three splits and $c=0.5$. Analogously, Figure 15 and Table 4 give the results for the setting with $c=2$.
Similar to the results in the settings with two splits, the estimated variance was close to the true variance.
Table 3: Ratio of the true standard error to the bootstrap estimated standard error of the debiased ECC estimators in the setting with three splits and $c=0.5$. The entries show summary statistics of the ratios calculated across the 100 different $\lambda$ values. Ratios greater than 1 indicate that the true standard error is greater than the bootstrap estimated standard error.
Estimator Min 1st Qu. Median Mean 3rd Qu. Max
Integral-based estimator 1.026 1.027 1.031 1.031 1.034 1.042
Newey-Robins estimator 1.014 1.016 1.024 1.023 1.029 1.031
Doubly robust estimator 0.990 1.015 1.016 1.016 1.018 1.019
Figure 14: Bootstrap estimated (black dots) versus true (red line) variance of the debiased ECC estimators in the setting with three splits and $c=0.5$ across the $\lambda$ values.
Table 4: Ratio of the true standard error to the bootstrap estimated standard error of the debiased ECC estimators in the setting with three splits and $c=2$. The entries show summary statistics of the ratios calculated across the 100 different $\lambda$ values. Ratios greater than 1 indicate that the true standard error is greater than the bootstrap estimated standard error.
Estimator Min 1st Qu. Median Mean 3rd Qu. Max
INT (integral-based) 0.934 0.935 0.939 0.941 0.944 0.978
NR (Newey–Robins) 0.935 0.935 0.940 0.942 0.946 0.977
IF (doubly robust) 0.935 0.936 0.943 0.946 0.951 0.994
Figure 15: Bootstrap estimated (black dots) versus true (red line) variance of the debiased ECC estimators in the setting with three splits and $c=2$ across the $\lambda$ values.